# failed: manyfoot

### v1

@@ 0.4 Conclusions
We have demonstrated a site-defined control architecture using foundry microelectronics, for the scalable manipulation of spin qubits. With NVs lithographically positioned at predefined sites, we showed that their associated electron spin can be driven efficiently, important for quantum computing, which requires liquid helium operation to extract coherent photons from NVs, and hence has limited thermal budget. Most importantly, our results demonstrate that crosstalk in driving adjacent spins can be strongly mitigated through an active cancellation scheme, without significantly sacrificing this driving efficiency. The scheme works without frequency detuning spin qubit transitions [25, 24, 26], allowing distinct control of sets of identical qubits.
The extinction ratio of field cancellation can also be improved by precisely balancing the driving of the inner and outer loops (see Methods).
@@
Foundry microelectronics also gives access to mature silicon device libraries. In combination with photonic integrated circuits, a large-scale spin-photon interfaced quantum information processor could be realised, overcoming space and energy constraints. In Fig.5(a), we present an exemplar multi-purpose device combining CMOS-compatible integrated photonics and multi-layer electronics for solid-state spin systems. We show in Fig.5(b) that besides spin driving, one could realise static field tuning [27], electrical spin readout [28, 29], photon detection, and logical electronics [30] in the same platform. The loop structure we present in the work (Fig.2(a)), with the central 10 µm region, is compatible with out-of-plane photonics (Fig.5(c-d)), such as bullseye gratings [31] or solid immersion lenses [32]. With the exit, our structures can also be combined with in-plane photonics (Fig.5(e-f)) such as waveguides [33] and photonic crystal nanocavities [34, 35]. Photonic accessibility provided by this scheme would greatly enhance the single photon collection efficiency and the coherent extraction of single photons.
Our demonstration consists of spin sites separated by , limited by the 0.13µm process node. However, it is possible to pursue higher densities at smaller process nodes by exploiting the presented crosstalk cancellation scheme. Considering demonstrated NV positioning accuracy of several hundreds nanometers [33] and a sub-micron optical modes through confocal microscopy or waveguide mode excitation [36], separation should be achievable. The scheme presented here is also translatable to alternative quantum platforms that use microwave driving, including other colour centres in diamond [37], single spin molecules [38], quantum dots [39] or superconducting circuits [40].
0.5 Methods
Device fabrication
We employ the IHP-SG13G2 BiCMOS technology with two thick metal layers for current-carrying applications and five thin metal layers for interconnections. The track width of the inner (outer) loops are (), both on the same thick aluminium layer. We follow the lithography method in Ref. [33] for nanodiamond positioning on the chip. Nanodiamonds are sized around diameter, milled from high pressure high temperature diamond (Nabond). First, a post-process inductively coupled plasma etch is used to release the inner region of the device, creating a -deep trench. This allows nanodiamonds to be positioned at the centre of the loops. A layer of polymethyl methacrylate (PMMA) is then coated and patterned by electron beam lithography (Raith Voyager). This create sites on the PMMA mask on which the nanodiamond (suspended in methanol) is then dropcast. Removing the PMMA then results in nanodiamonds positioned precisely on the predefined sites.
Experimental setup
A scanning confocal microscope is used to optically address the NV in this work. With a 0.9 numerical aperture microscope objective (Olympus M Plan N 100x/0.90NA), the excitation beam (515 nm, Cobolt 06-MLD) is focused on the sample at a nearly diffraction-limited spot ( diameter). The NV photoluminescence is collected through the same lens and separated from the excitation path using a dichroic mirror. By using a high NA objective, the NV PL is extracted from the background fluorescence such as scattering from metal layers. Electronically, for the active cancelling demonstration, the microwave source (Rohde & Schwarz SMB100AP20) is split by a power splitter (Mini-Circuits ZN2PD2-63-S+) for driving the inner and outer right metallic loop interface. On one path, an additional tunable power attenuator (Mini-Circuits RCDAT-6G-120H) and phase shifter (Vaunix LPS-802) are applied. Both paths are amplified (Microwave Amplifiers Ltd AM4-2-6-43-43R) and fed to an isolator (DiTom Microwave D3C2040) before the device-under-test. Timetagging and pulsing are recorded by UQDevices Logic-16 and Swabian Instruments Pulse Streamer 8/2. The electronic setup for crosstalk-mitigated spin control only differs with another random phase shifter applied before power splitting. Pulsing is achieved using a microwave switch (Mini-Circuits, ZASW-2-50DR+, DC-5 GHz) in Rabi oscillation measurements. Full electronic circuits are shown in Supplementary Fig.S6.
Electron spin Rabi oscillations and CW-ODMR measurements
We perform electron spin Rabi oscillations by driving the NV ground state triplets ( and ). Fig.3a shows the corresponding pulse sequences. The electron spin is first polarised into the state by a long 515 nm laser pulse. A microwave pulse of length is then applied to resonantly drive the transition between (with the ground state zero-field splitting around 2.87 GHz). In our measurement, an external static magnetic field is applied using a strong permanent magnet in order to lift the degeneracy of the two transitions ( and ) and to shift the resonances such that it is best matched to the on-chip loop resonance (Supplementary Section E). Another laser pulse of is used to readout the spin state by the PL intensity (stronger in the state and weaker in the state due to state-dependent intersystem crossing). The PL intensity is extracted by summing the photon count in the first interval and normalising to the count summed over the last . We use the same laser pulse to polarise and readout the electron state. In Fig.3b-j, the Rabi measurements are presented (in the y-axis) with the relative intensity normalised between for comparison. For the Rabi oscillation results presented in Fig.4h-k, the error in decay time is deduced from the variance of the parameter estimation during fitting, presenting one standard deviation error. The slight drift in oscillation is an artifact due to device heating under microwave power.
In the continuous-wave ODMR measurement, the 515 nm laser is constantly on while the frequency of the microwave source is swept around the resonance of electron spin transitions. Again, due to the difference of PL intensity for the electron states, a dip is observed at resonances with a certain depth (the ODMR contrast). For results shown in Fig.4a-e, the contrast is extracted by fitting the measured PL intensity across frequency sweeps to a Gaussian dip. Error bars, showing one standard deviation in photon counts, arise from the photon shot noise and low drift of the experimental setup.
Extinction ratio of crosstalk field
The field cancellation extinction ratio (directly proportional to the residual crosstalk field) is undetectable (within experimental error bar) from the CW-ODMR measurements (Fig.4c-e) and Rabi oscillation lifetimes (Fig.4h-k) directly. Fitting the results, a -15 dB extinction ratio of crosstalk field power is inferred. The extinction ratio can be further improved by precisely balancing the driving of the inner-outer loops to the optimal ratio. For our case demonstrated in Fig.4(h-k), a -20 dB imbalance of microwave power (from the left loops) at the NV would diminish only by 2% under the active cancelling protection and a -30 dB imbalance for 0.2% (Supplementary Section I).
[…]
Acknowledgements
@@
The authors thank J. Monroy-Ruz for building the confocal microscope used in this experiment and S. Komrakova and J. Frazer for fabrication support. The authors thank J. Monroy-Ruz and J. Matthews for useful discussions. The authors acknowledge funding support from the Quantum Computing and Simulation Hub Partnership Resource Project RFSQ and the Engineering and Physical Sciences Research Council (EPSRC) grant QC:SCALE EP/W006685/1.

### v2

@@ V Conclusions
We have demonstrated a site-defined control architecture using foundry microelectronics, for the scalable manipulation of spin qubits. With NVs lithographically positioned at predefined sites, we showed that their associated electron spin can be driven efficiently, important for quantum computing, which requires liquid helium operation to extract coherent photons from NVs, and hence has limited thermal budget. Most importantly, our results demonstrate that crosstalk in driving adjacent spins can be strongly mitigated through an active cancellation scheme, without significantly sacrificing this driving efficiency. The scheme works without frequency detuning spin qubit transitions [28, 27, 29], allowing distinct control of sets of identical qubits. Note that our crosstalk mitigation method through hardware design prevents the time-consuming pair-wise calibration and sophisticated signal compensation typically used in addressing qubit systems [30]. The extinction ratio of the field cancellation can also be improved by more precisely balancing the driving of the inner and outer loops (see Appendix D). In this work, nanodiamonds are primarily used for ease of prototyping. For large scale quantum information applications, use of implanted NVs in strain-minimised diamond membrane structure [31] would minimise difference in local NV properties which could also be locally tuned [32].
Fig. 5 showcases how the double-loop crosstalk-mitigated microelectronics would fit into existing large-scale quantum information processing hardware for optically active spins [33, 34, 35]. The micron-scale separation of spin defects in the architecture highlights the need of crosstalk-mitigation microelectronics. Fig. 5a shows the integration of a 16 by 16 spin array with input and output photonic links for selective addressing and entangling photonic gates. Notably the photonic resonant transition of NVs can be tuned electrically [34], which allows single spins along the waveguide to be selected and addressed separately [36]. The tunability also enables multi-qubit gate or entanglement of spins across waveguides through interference of emitted photons [37] of the same wavelength. The scheme could be readily adopted to previous work [38, 35, 34]. Technological challenges remain in realising a platform that combines a visible wavelength photonics with multilayer microelectronics, although this has seen dramatic progress in recent years with silicon nitride integration [39, 40], and the subsequent integration of the thin film diamond containing the qubits [35]. With qubit spacing of 15 µm and the inner(outer) loop dimension of 3(6) µm, the whole hundred-physical-qubit device has a dimension 1mm2, demonstrating high level of integration. It is also possible to pursue higher densities at smaller process nodes by exploiting the presented crosstalk cancellation scheme and significant headroom in the current capacity of the existing circuit. Considering NV positioning accuracy of 50 nm and a sub-micron optical modes through confocal microscopy or waveguide mode excitation [41], sub-micron separation should be achievable.
@@
Foundry microelectronics also gives access to mature silicon device libraries to realise static field tuning [42], electrical spin readout [43, 44], photon detection, and logical electronics [20] in the same platform. In combination with photonic integrated circuits, a large-scale spin-photon interfaced quantum information processor could be realised, overcoming space and energy constraints. The scheme presented here is also translatable to alternative quantum platforms that use microwave driving, including other colour centres in diamond [45], single spin molecules [46], quantum dots [47] or superconducting circuits [48].
Acknowledgements.
@@
The authors thank J. Monroy-Ruz for building the confocal microscope used in this experiment and S. Komrakova, J. Frazer, and F. Malik for fabrication support. The authors thank J. Monroy-Ruz and J. Matthews for useful discussions. The authors acknowledge funding support from the Quantum Computing and Simulation Hub Partnership Resource Project RFSQ and the Engineering and Physical Sciences Research Council (EPSRC) grant QC:SCALE EP/W006685/1.Appendix A Analytical model expression
The analytical model presented (Fig. 1 in the main text) uses the static Biot-Savart approximation. We consider the off-axis field of a current loop centred at the origin in the x-y plane. For any point in the space, the magnetic field is defined in the cylindrical coordinate system with the component in axial direction
and in the radial direction
where and . Here, is the permeability constant, is the radius of loop, is the current flow in the loop wire, is the elliptic integral function of first kind, is the elliptic integral function of second kind, and constants defined as , , , and .
Appendix B Finite-element simulation details
To verify the device design, a finite-element model of the structure is carried out in COMSOL Multiphysics following the analytical model. The double-loop structure is constructed with aluminium metallic loops surrounded by silicon oxide with relative permittivity . Lumped ports are defined between the input and output metallic wire of the loops and excited by a voltage source with the port phase set in simulation. The Electromagnetic wave, Frequency domain analysis COMSOL package is used to study the device at 2.87 GHz. The minor difference between analytical and simulation results is due to the physical width of the wires of loop structure that is not considered in the analytical model.
Appendix C Device fabrication details
We employ the IHP-SG13G2 BiCMOS technology with two thick metal layers for current-carrying applications and five thin metal layers for interconnections. The track width of the inner (outer) loops are (), both on the same thick aluminium layer.
For nanodiamond positioning, nanodiamonds are sized around diameter, milled from high pressure high temperature diamond (Nabond). First, a post-process inductively coupled plasma etch is used to release the inner region of the device, creating a -deep trench. This allows nanodiamonds to be positioned at the centre of the loops. A layer of polymethyl methacrylate (PMMA) is then coated and patterned by electron beam lithography (Raith Voyager). This creates sites on the PMMA mask on which the nanodiamond (suspended in methanol) is then dropcast. Removing the PMMA then results in nanodiamonds positioned precisely on the predefined sites.
Appendix D Experimental setup
A scanning confocal microscope is used to optically address the NV in this work. With a 0.9 numerical aperture microscope objective (Olympus M Plan N 100x/0.90NA), the excitation beam (515 nm, Cobolt 06-MLD) is focused on the sample at a nearly diffraction-limited spot ( diameter). The NV photoluminescence is collected through the same lens and separated from the excitation path using a dichroic mirror. By using a high NA objective, the NV PL is extracted from the background fluorescence such as scattering from metal layers. Electronically, for the active cancelling demonstration, the microwave source (Rohde & Schwarz SMB100AP20) is split by a power splitter (Mini-Circuits ZN2PD2-63-S+) for driving the inner and outer right metallic loop interface. On one path, an additional tunable power attenuator (Mini-Circuits RCDAT-6G-120H) and phase shifter (Vaunix LPS-802) are applied. Both paths are amplified (Microwave Amplifiers Ltd AM4-2-6-43-43R) and fed to an isolator (DiTom Microwave D3C2040) before the device-under-test. Timetagging and pulsing are recorded by UQDevices Logic-16 and Swabian Instruments Pulse Streamer 8/2. The electronic setup for crosstalk-mitigated spin control only differs with another random phase shifter applied before power splitting. Pulsing is achieved using a microwave switch (Mini-Circuits, ZASW-2-50DR+, DC-5 GHz) in Rabi oscillation measurements.
We perform electron spin Rabi oscillations by driving the NV ground state triplets ( and ). Fig. 3a shows the corresponding pulse sequences. The electron spin is first polarised into the state by a long 515 nm laser pulse. A microwave pulse of length is then applied to resonantly drive the transition between (with the ground state zero-field splitting around 2.87 GHz). In our measurement, an external static magnetic field is applied using a strong permanent magnet in order to lift the degeneracy of the two transitions ( and ) and to shift the resonances such that it is best matched to the on-chip loop resonance. Another laser pulse of is used to readout the spin state by the PL intensity (stronger in the state and weaker in the state due to state-dependent intersystem crossing). The PL intensity is extracted by summing the photon count in the first interval and normalising to the count summed over the last . We use the same laser pulse to polarise and readout the electron state. In Fig. 3b-g, the Rabi measurements are presented (in the y-axis) with the relative intensity normalised between for comparison. For the Rabi oscillation results presented in Fig. 4e-h, the error in decay time is deduced from the variance of the parameter estimation during fitting, presenting one standard deviation error. The slight drift in oscillation is an artifact due to device heating under microwave power.
In the continuous-wave ODMR measurement, the 515 nm laser is constantly on while the frequency of the microwave source is swept around the resonance of electron spin transitions. Again, due to the difference of PL intensity for the electron states, a dip is observed at resonances with a certain depth (the ODMR contrast). The contrast is extracted by fitting the measured PL intensity across frequency sweeps to a Gaussian dip. Error bars, showing one standard deviation in photon counts, arise from the photon shot noise and low drift of the experimental setup.
The field cancellation extinction ratio (directly proportional to the residual crosstalk field) is undetectable (within experimental error bar) from the CW-ODMR measurements and Rabi oscillation lifetimes (Fig. 4e-h) directly. Fitting the results in Fig. 4a, a -15 dB extinction ratio of crosstalk field power is inferred. The extinction ratio can be further improved by precisely balancing the driving of the inner-outer loops to the optimal ratio. For our case demonstrated in Fig. 4(e-h), a -20 dB imbalance of microwave power (from the left loops) at the NV would diminish only by 2% under the active cancelling protection and a -30 dB imbalance for 0.2%.